An assembly operation is to be done in a single shift, 8 hours per day and 10% allowed for breaks (PF&D). The operators are expected to have an 85% performance and the following parts are required:

Part X, requirement is 10,000 per week, at 100% one operator can produce 150 pieces per hour.

Part Y, requirement is 15,000 per week, at 100% one operator can produce 220 pieces per hour.

Part Z, requirement is 20,000 per week, at 100% one operator can produce 275 pieces per hour.

QUESTIONS TO ANSWER:

3. Assuming no setup time, how many people are needed to meet production requirements?

4. How many people are needed if performance is expected to be only 75%?
Answer #1

IF YOU LIKE THE ANSWER, PLEASE GIVE AN UP-VOTE OR THUMB UP. THIS WILL ENCOURAGE ME TO ANSWER MORE!!

Shift is 8 hours a day. Means 480 minutes a day. 10% allowances for PF&D. Hence working minutes a day = 480-48 = 432 minutes

Hence working hours per week = 432*7 = 3024 minutes

For Part X

Requirement is 10000 a week

Performance is 85%

At 100% labor can produce 150 pieces per hour. At 85%, labor will produce 127.5 parts an hour. Hence in 60 minutes 127.5 parts

In a week one labor can produce: (127.5/60)*3024 = 6426 parts

Labor weeks needed for 10000 parts a week = 10000/6426 = 1.56 labor weeks

For Part Y

Requirement is 15000 a week

Performance is 85%

At 100% labor can produce 220 pieces per hour. At 85%, labor will produce 187 parts an hour. Hence in 60 minutes 187 parts

In a week one labor can produce: (187/60)*3024 = 9424.8 parts

Labor weeks needed for 15000 parts a week = 15000/9424.8 = 1.66 labor weeks

For Part Z

Requirement is 20000 a week

Performance is 85%

At 100% labor can produce 275 pieces per hour. At 85%, labor will produce 233.75 parts an hour. Hence in 60 minutes 233.75 parts

In a week one labor can produce: (233.75/60)*3024 = 11781 parts

Labor weeks needed for 15000 parts a week = 20000/11781 = 1.69 labor weeks

Hence total labors required = 1.56+1.66+1.69 = 4.91 = 5 labors

b.

For Part X

Requirement is 10000 a week

Performance is 75%

At 100% labor can produce 150 pieces per hour. At 75%, labor will produce 112.5 parts an hour. Hence in 60 minutes 112.5 parts

In a week one labor can produce: (112.5/60)*3024 = 5670 parts

Labor weeks needed for 10000 parts a week = 10000/5670 = 1.76 labor weeks

For Part Y

Requirement is 15000 a week

Performance is 75%

At 100% labor can produce 220 pieces per hour. At 75%, labor will produce 165 parts an hour. Hence in 60 minutes 165 parts

In a week one labor can produce: (165/60)*3024 = 8316 parts

Labor weeks needed for 15000 parts a week = 15000/8316 = 1.80 labor weeks

For Part Z

Requirement is 20000 a week

Performance is 75%

At 100% labor can produce 275 pieces per hour. At 75%, labor will produce 206.25 parts an hour. Hence in 60 minutes 206.25 parts

In a week one labor can produce: (206.25/60)*3024 = 10395 parts

Labor weeks needed for 15000 parts a week = 20000/10395 = 1.92 labor weeks

Hence total labors required = 1.76+1.80+1.92 = 5.48 = 6 labors